### v1

@@ A.6 Setting of $K$ for (32)
A.7 Proof of Corollary 4.5
The binomial distribution (33) belongs to the exponential family (11) with $\bm{\gamma}(\bm{x})=\bm{x}$, so it is easy to see that for $\bm{\Omega}=\mathrm{diag}(\frac{1}{m_{1}},\ldots,\frac{1}{m_{p}})$, we have $\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}[\bm{\Omega}\bm{\gamma}(\bm{X})]=\bm{\theta}$. In addition, since $\bm{\eta}(\bm{\theta})=(\ln{\frac{\theta_{1}}{1-\theta_{1}}},\ldots,\ln{\frac{\theta_{p}}{1-\theta_{p}}})$, we have $\bm{\Omega}\big{(}\nabla\bm{\eta}(\bm{\theta})\big{)}^{-1}=\mathrm{diag}(\frac{\theta_{1}(1-\theta_{1})}{m_{1}},\ldots,\frac{\theta_{p}(1-\theta_{p})}{m_{p}})$, which is equal to $\mathbb{V}_{p_{\bm{x}}(\bm{\theta})}[\bm{\Omega}\bm{\gamma}(\bm{X})]=\mathrm{diag}(\frac{\theta_{1}(1-\theta_{1})}{m_{1}},\ldots,\frac{\theta_{p}(1-\theta_{p})}{m_{p}})\equiv\bm{I}_{p_{\bm{x}}}(\bm{\theta})^{-1}$. Substituting this into (19), we get $\bm{\theta}^{(t)}\leftarrow\bm{\theta}^{(t-1)}-\mathrm{diag}\Big{(}\frac{\theta^{(t-1)}_{1}(1-\theta^{(t-1)}_{1})}{m_{1}},...,\frac{\theta^{(t-1)}_{p}(1-\theta^{(t-1)}_{p})}{m_{p}}\Big{)}\nabla f(\bm{\theta}^{(t-1)}).$
A.8 Proof of Proposition 4.6
@@
$\mathbb{E}_{p_{x}(\theta)}\Big{[}\prod_{j\in[p]}X_{j}(X_{j}-1)\cdots(X_{j}-n_{j}+1)\Big{]}$ $=m(m-1)\cdots\Big{(}m-\sum_{j\in[p]}n_{j}+1\Big{)}\theta_{1}^{n_{1}}\cdots\theta_{p}^{n_{p}},$ (78)
where $X_{p}=m-\sum\limits_{j\in[p-1]}X_{j},\theta_{p}=1-\sum\limits_{j\in[p-1]}\theta_{j}$. The formula (78) holds since we have
@@
where the last equality holds because the last term of the second line from the bottom is the summation of the probability mass function $p_{\bm{x}^{\prime}}(\bm{\theta})$ of $\mathrm{Multi}(m-\sum\limits_{j\in[p]}n_{j},\bm{\theta})$. (78) implies that $\mathbb{E}_{p_{x}(\theta)}\big{[}g_{i}(\bm{X})\big{]}=\prod\limits_{j\in[p]}\theta_{j}^{n_{i,j}}$ and that
@@
$\equiv K-F(\bm{\theta}).$
A.9 Setting of $K$ for (41)
Similar to Section A.6, this subsection shows how the value $K$ can be given for (41). Since $x_{j}\in\{0,1,\ldots,m\}$ and $m\geq\sum\limits_{j\in[p]}n_{i,j}>0$, we have $\prod\limits_{j\in[p]}\prod\limits_{k\in[n_{i,j}]}(x_{j}-k+1)\leq\prod\limits_{k\in[\sum_{j\in[p]}n_{i,j}]}(m-k+1)$ and thus $g_{i}(\bm{x})\leq 1$ by the definition of $g_{i}$. This leads to $a_{i}g_{i}(\bm{x})\leq a_{i}$ for $a_{i}\geq 0$; $a_{i}g_{i}(\bm{x})\leq 0$ for $a_{i}<0$, and $G(\bm{x}):=\sum\limits_{i\in[I]}a_{i}\prod\limits_{j\in[p]}g_{i,j}(x_{j})\leq\sum\limits_{i\in[I]}\max\{a_{i},0\},$ so the constant defined as $K:=\sum\limits_{i\in[I]}\max\{a_{i},0\}$ gives the upper bound of the optimal value.
A.10 Proof of Corollary 4.7
With $\bm{\gamma}(\bm{X})=\bm{X},\bm{\Omega}=\frac{1}{m}\bm{I}\in\mathbb{R}^{(p-1)\times(p-1)}$, $\bm{\eta}(\bm{\theta})=(\ln\frac{\theta_{1}}{1-\bm{1}^{\top}\bm{\theta}},\ldots,\ln\frac{\theta_{p-1}}{1-\bm{1}^{\top}\bm{\theta}})^{\top}$, (17) is valid. The $(k,l)$-elements of the inverse of the Fisher information matrix is
$\big{(}\bm{I}_{p_{\bm{x}}}(\bm{\theta})^{-1}\big{)}_{k,l}\equiv\big{(}\mathbb{V}_{p_{\bm{x}}(\bm{\theta})}[\bm{\Omega}\bm{\gamma}(\bm{X})]\big{)}_{k,l}=\begin{cases}\frac{\theta_{k}(1-\theta_{k})}{m},&k=l\land k,l\in[p-1],\\ -\frac{\theta_{k}\theta_{l}}{m},&k\neq l\land k,l\in[p-1],\end{cases}$
which follow from (18) and
@@ A.9 Setting of $K$ for (41)
we get the update formula (43).
A.11 Proof of Proposition 5.2
@@
A.12 Proof of Theorem 4.8
The optimality condition for (16) implies
$\nabla f(\bm{\theta}^{(t)})+\nabla\bm{\eta}(\bm{\theta}^{(t)})^{\top}\big{(}\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t)})}[\bm{\gamma}(\bm{X})]-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]\big{)}$ $=\bm{0}.$ (79)
By the definition (47) of $\bm{\eta}(\bm{\theta})$, the second term on the left-hand side of (79) is
where we denote $\bm{\mu}(\bm{\theta})=(\mu_{1}(\bm{\theta}),\ldots,\mu_{p}(\bm{\theta}))^{\top}:=\mathbb{E}_{q_{\bm{x}}(\bm{\theta})}[\bm{\gamma}(\bm{X})]$. With $\nu_{i}^{(t)}:=\sum_{j\in[p]}r_{j,i}(\mu_{j}(\bm{\theta}^{(t)})-\mu_{j}(\bm{\theta}^{(t-1)}))/g_{i}(\bm{\theta}^{(t)})$, we have $\nu_{i}^{(t)}g_{i}(\bm{\theta}^{(t)})=\sum_{j\in[p]}r_{j,i}(\mu_{j}(\bm{\theta}^{(t)})-\mu_{j}(\bm{\theta}^{(t-1)}))\to 0$ as $t\to\infty$, and (79) is represented as
$\nabla f(\bm{\theta}^{(t)})+\sum\limits_{i\in[m]}\nu_{i}^{(t)}\nabla g_{i}(\bm{\theta}^{(t)})$ $=\bm{0}.$ (80)
If (48) is satisfied, we can let $\nu^{*}_{i}:=\lim\limits_{t\to\infty}\nu^{(t)}_{i}$ and see that the KKT condition is fulfilled at the limit $\bm{\theta}=\bm{\theta}^{*}$: $\nabla f(\bm{\theta}^{*})+\sum_{i\in[m]}\nu_{i}^{*}\nabla g_{i}(\bm{\theta}^{*})=\bm{0},~{}g_{i}(\bm{\theta}^{*})\leq 0,~{}\nu_{i}^{*}g_{i}(\bm{\theta}^{*})=0,~{}\nu_{i}^{*}\geq 0,~{}i\in[m].$
A.13 Fulfillment of the condition (48) of Theorem 4.8
Case (i): Algorithm 3
The box constraints of (32) can be represented as in the form of (46) with $g_{j}(\bm{\theta})=-\theta_{j},~{}g_{p+j}(\bm{\theta})=\theta_{j}-1,~{}j\in[p]$ (i.e., $m=2p$). It is easy to see that the parameters $\bm{\eta}(\bm{\theta})=\big{(}\ln\frac{\theta_{1}}{1-\theta_{1}},\ldots,\ln\frac{\theta_{p}}{1-\theta_{p}}\big{)}^{\top}$ for the binomial distribution (33) are represented as (47) with $r_{j,i}:=1~{}(i=j);~{}-1~{}(i=j+p);~{}0$ (otherwise) for $i\in[2p],~{}j\in[p]$. The formula (15) then implies that $\mu_{j}(\bm{\theta}^{(t-1)})=m_{j}\theta_{j}^{(t-1)}-\theta_{j}^{(t-1)}(1-\theta_{j}^{(t-1)})\frac{\partial}{\partial\theta_{j}}f(\bm{\theta}^{(t-1)}),j\in[p].$ To see (48), observe that for the $i$-th constraint $g_{i}(\bm{\theta})=-\theta_{i},~{}r_{i,i}=1,~{}i\in[p]$, we have
which converges to $\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{*})\geq 0$ as $\theta^{(t)}_{i}\to 0$ because it would contradict to the monotonicity of $\{f(\bm{\theta}^{(t)})\}$ otherwise. The fourth equality follows from the update formula (35), which implies that $m_{i}\big{(}\theta^{(t)}_{i}-\theta^{(t-1)}_{i}\big{)}=\theta^{(t)}_{i}(1-\theta^{(t-1)}_{i})\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{(t-1)})$ Similarly, for the $(p+i)$-th constraint $g_{i+p}(\bm{\theta})=\theta_{i}-1,~{}r_{i+p,i+p}=-1,~{}i\in[p]$, we have $\nu^{(t)}_{i}=-\theta^{(t)}_{i}\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{(t)}),$ which converges to $-\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{*})\geq 0$ as $\theta^{(t)}_{i}\to 1$ because it would contradict to the monotonicity of $\{f(\bm{\theta}^{(t)})\}$ otherwise. Consequently, in either constraint, we can see that $\lim\limits_{t\to\infty}\nu_{i}^{(t)}\geq 0$.
Case (ii): Algorithm 4
The simplex constraint of (41) corresponds to $g_{j}(\bm{\theta})=-\theta_{j},~{}j\in[p-1]$, and $g_{p}(\bm{\theta})=\bm{1}^{\top}\bm{\theta}-1$. With $r_{j,i}:=1~{}(i=j);~{}-1~{}(i=p);~{}0$ (otherwise) for $i\in[p],~{}j\in[p-1]$, the parameters $\bm{\eta}(\bm{\theta})=\big{(}\ln\frac{\theta_{1}}{1-\bm{1}^{\top}\bm{\theta}},\ldots,\ln\frac{\theta_{p-1}}{1-\bm{1}^{\top}\bm{\theta}}\big{)}^{\top}$ for the multinomial distribution are represented in the form of (47). The formula (15) then implies that
$\mu_{j}(\bm{\theta}^{(t-1)})=m_{j}\theta_{j}^{(t-1)}-\big{(}\theta_{j}^{(t-1)}\frac{\partial}{\partial\theta_{j}}f(\bm{\theta}^{(t-1)})-\theta_{j}^{(t-1)}(\bm{\theta}^{(t-1)})^{\top}\nabla f(\bm{\theta}^{(t-1)})\big{)},~{}j\in[p-1].$
To see (48), observe that for the $i$-th constraint $g_{i}(\bm{\theta})=-\theta_{i},~{}r_{i,i}=1,~{}i\in[p-1]$, we have
$\nu^{(t)}_{i}=\frac{\mu_{i}(\bm{\theta}^{(t)})-\mu_{i}(\bm{\theta}^{(t-1)})}{-\theta_{i}^{(t)}}=(1-\theta_{i}^{(t)})\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{(t)})-\sum_{j\neq i}\theta_{j}^{(t)}\frac{\partial}{\partial\theta_{j}}f(\bm{\theta}^{(t)})$ (81)
The second equality follows from the update formula (43), which implies that
$m_{i}\big{(}\theta^{(t)}_{i}-\theta^{(t-1)}_{i}\big{)}=-\theta_{i}^{(t-1)}\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{(t-1)})+\theta_{i}^{(t-1)}(\bm{\theta}^{(t-1)})^{\top}\nabla f(\bm{\theta}^{(t-1)}).$
We see from (81) that $\nu^{(t)}_{i}$ converges to a finite value $\nu^{*}_{i}$. To see the sign of $\nu^{*}$, observe that if $\lim\limits_{t\to\infty}\theta^{(t)}_{i}=0$, the first term of (81) is nonnegative because $\frac{\partial}{\partial\theta_{i}}f(\bm{\theta})$ is continuous and the optimality condition implies $\lim\limits_{t\to\infty}\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{(t)})\geq 0$. We can also show $\lim\limits_{t\to\infty}\theta^{(t)}_{i}=1$ similarly. As for the case where $\lim\limits_{t\to\infty}\theta^{(t)}_{i}\in(0,1)$, it converges to zero since $\lim\limits_{t\to\infty}\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{(t)})=0$ must hold then. As for the second term of (81), we see that $\lim\limits_{t\to\infty}-\theta_{j}^{(t)}\frac{\partial}{\partial\theta_{j}}f(\bm{\theta}^{(t)})\geq 0$ for every $j\neq i$, likewise. Similarly, for the $p$-th constraint $g_{p}(\bm{\theta})=\bm{1}^{\top}\bm{\theta}-1$, we have
A.14 Proof of Proposition 5.3
@@
Let $\phi(t)=t\ln t-t+1$ and $H(\bm{x},\bm{y}):=\sum_{j\in[p]}y_{j}\phi(\frac{x_{j}}{y_{j}})\equiv x_{j}\ln\frac{x_{j}}{y_{j}}+y_{j}-x_{j}$, and $L(\bm{\theta}):=(l_{1}(\bm{\theta}),...,l_{p}(\bm{\theta}))$ where $l_{j}(\bm{\theta}):=c_{j}-\bm{a}_{j}^{\top}\bm{\theta},j\in[p]$. Here, $H(\bm{x},\bm{y})$ denotes the KL divergence for $\bm{x}\in\mathbb{R}^{p}_{+}$ relative to $\bm{y}\in\mathbb{R}^{p}_{++}$. Observe that Algorithm 7 is a proximal-like algorithm which iteratively solves
@@
By substituting $\bm{\theta}=\bm{\theta}^{*}$, the right-hand side goes to 00 as $t\to\infty$. The second statement can be shown in the same manner as Theorem 4.3(iii) of [31] by extending $H(L(\bm{\theta}),L(\bm{\theta}^{(t)}))$ to the composite proximal term, $H(L(\bm{\theta}),L(\bm{\theta}^{(t)}))+\frac{1}{2}\|\bm{\theta}-\bm{\theta}^{(t)}\|_{\bm{\Sigma}^{-1}-\bm{Q}}^{2}$.

### v2

@@ Appendix A Derivation of formulae and proof of propositions
A.7. Proof of Corollary 4.5
The binomial distribution (4.17) belongs to the exponential family (3.1) with $\bm{\gamma}(\bm{x})=\bm{x}$, so it is easy to see that for $\bm{\Omega}=\mathrm{diag}(\frac{1}{m_{1}},\ldots,\frac{1}{m_{p}})$, we have $\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}[\bm{\Omega}\bm{\gamma}(\bm{X})]=\bm{\theta}$. In addition, since $\bm{\eta}(\bm{\theta})=(\ln{\frac{\theta_{1}}{1-\theta_{1}}},\ldots,\ln{\frac{\theta_{p}}{1-\theta_{p}}})$, we have $\bm{\Omega}\big(\nabla\bm{\eta}(\bm{\theta})\big)^{-1}=\mathrm{diag}(\frac{\theta_{1}(1-\theta_{1})}{m_{1}},\ldots,\frac{\theta_{p}(1-\theta_{p})}{m_{p}})$, which is equal to $\mathbb{V}_{p_{\bm{x}}(\bm{\theta})}[\bm{\Omega}\bm{\gamma}(\bm{X})]=\mathrm{diag}(\frac{\theta_{1}(1-\theta_{1})}{m_{1}},\ldots,\frac{\theta_{p}(1-\theta_{p})}{m_{p}})=\bm{I}_{p_{\bm{x}}}(\bm{\theta})^{-1}$. Substituting this into (3.9), we get $\bm{\theta}^{(t)}\leftarrow\bm{\theta}^{(t-1)}-\mathrm{diag}\Big(\frac{\theta^{(t-1)}_{1}(1-\theta^{(t-1)}_{1})}{m_{1}},...,\frac{\theta^{(t-1)}_{p}(1-\theta^{(t-1)}_{p})}{m_{p}}\Big)\nabla f(\bm{\theta}^{(t-1)}).$
A.8. Proof of Proposition 4.6
@@
$\mathbb{E}_{p_{x}(\theta)}\Big[\prod_{j\in[p]}X_{j}(X_{j}-1)\cdots(X_{j}-n_{j}+1)\Big]$ $=m(m-1)\cdots\Big(m-\sum_{j\in[p]}n_{j}+1\Big)\theta_{1}^{n_{1}}\cdots\theta_{p}^{n_{p}},$ (A.7)
where $X_{p}=m-\sum\limits_{j\in[p-1]}X_{j},\theta_{p}=1-\sum\limits_{j\in[p-1]}\theta_{j}$. The formula (A.7) holds since we have
@@
[…]
@@
$=K-F(\bm{\theta}).$
A.9. Setting of $K$ for (4.28)
Similar to Section A.6, this subsection shows how the value $K$ can be given for (4.28). Since $x_{j}\in\{0,1,\ldots,m\}$ and $m\geq\sum\limits_{j\in[p]}n_{i,j}>0$, we have $\prod\limits_{j\in[p]}\prod\limits_{k\in[n_{i,j}]}(x_{j}-k+1)\leq\prod\limits_{k\in[\sum_{j\in[p]}n_{i,j}]}(m-k+1)$ and thus $g_{i}(\bm{x})\leq 1$ by the definition of $g_{i}$. This leads to $a_{i}g_{i}(\bm{x})\leq a_{i}$ for $a_{i}\geq 0$; $a_{i}g_{i}(\bm{x})\leq 0$ for $a_{i}<0$, and $G(\bm{x}):=\sum\limits_{i\in[I]}a_{i}\prod\limits_{j\in[p]}g_{i,j}(x_{j})\leq\sum\limits_{i\in[I]}\max\{a_{i},0\},$ so the constant defined as $K:=\sum\limits_{i\in[I]}\max\{a_{i},0\}$ gives the upper bound of the optimal value.
A.10. Proof of Corollary 4.7
With $\bm{\gamma}(\bm{x})=\bm{x},\bm{\Omega}=\frac{1}{m}\bm{I}\in\mathbb{R}^{(p-1)\times(p-1)}$, $\bm{\eta}(\bm{\theta})=(\ln\frac{\theta_{1}}{1-\bm{1}^{\top}\bm{\theta}},\ldots,\ln\frac{\theta_{p-1}}{1-\bm{1}^{\top}\bm{\theta}})^{\top}$, (3.7) is valid. The $(k,l)$-elements of the inverse of the Fisher information matrix is
$\big(\bm{I}_{p_{\bm{x}}}(\bm{\theta})^{-1}\big)_{k,l}=\big(\mathbb{V}_{p_{\bm{x}}(\bm{\theta})}[\bm{\Omega}\bm{\gamma}(\bm{X})]\big)_{k,l}=\begin{cases}\frac{\theta_{k}(1-\theta_{k})}{m},&k=l\land k,l\in[p-1],\\ -\frac{\theta_{k}\theta_{l}}{m},&k\neq l\land k,l\in[p-1],\end{cases}$
which follow from (3.8) and
@@
we get the update formula (4.30).
A.11. Proof of Theorem 4.8
The optimality condition for (3.6) implies
$\nabla f(\bm{\theta}^{(t)})+\nabla\bm{\eta}(\bm{\theta}^{(t)})^{\top}\big(\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t)})}[\bm{\gamma}(\bm{X})]-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]\big)$ $=\bm{0}.$ (A.8)
By the definition (4.34) of $\bm{\eta}(\bm{\theta})$, the second term on the left-hand side of (A.8) is
where we denote $\bm{\mu}(\bm{\theta})=(\mu_{1}(\bm{\theta}),\ldots,\mu_{p}(\bm{\theta}))^{\top}:=\mathbb{E}_{q_{\bm{x}}(\bm{\theta})}[\bm{\gamma}(\bm{X})]$. With $\nu_{i}^{(t)}:=\sum_{j\in[p]}r_{j,i}(\mu_{j}(\bm{\theta}^{(t)})-\mu_{j}(\bm{\theta}^{(t-1)}))/g_{i}(\bm{\theta}^{(t)})$, we have $\nu_{i}^{(t)}g_{i}(\bm{\theta}^{(t)})=\sum_{j\in[p]}r_{j,i}(\mu_{j}(\bm{\theta}^{(t)})-\mu_{j}(\bm{\theta}^{(t-1)}))\to 0$ as $t\to\infty$, and (A.8) is represented as
$\nabla f(\bm{\theta}^{(t)})+\sum\limits_{i\in[M]}\nu_{i}^{(t)}\nabla g_{i}(\bm{\theta}^{(t)})$ $=\bm{0}.$ (A.9)
If (4.35) is satisfied, we can let $\nu^{*}_{i}:=\lim\limits_{t\to\infty}\nu^{(t)}_{i}$ and see that the KKT condition is fulfilled at the limit $\bm{\theta}=\bm{\theta}^{*}$: $\nabla f(\bm{\theta}^{*})+\sum_{i\in[M]}\nu_{i}^{*}\nabla g_{i}(\bm{\theta}^{*})=\bm{0},~g_{i}(\bm{\theta}^{*})\leq 0,~\nu_{i}^{*}g_{i}(\bm{\theta}^{*})=0,~\nu_{i}^{*}\geq 0,~i\in[M].$
A.12. Fulfillment of the condition (4.35) of Theorem 4.8
Case (i): Algorithm 3
The box constraints of (4.16) can be represented as in the form of (4.33) with $g_{j}(\bm{\theta})=-\theta_{j},~g_{p+j}(\bm{\theta})=\theta_{j}-1,~j\in[p]$ (i.e., $M=2p$). It is easy to see that the parameters $\bm{\eta}(\bm{\theta})=\big(\ln\frac{\theta_{1}}{1-\theta_{1}},\ldots,\ln\frac{\theta_{p}}{1-\theta_{p}}\big)^{\top}$ for the binomial distribution (4.17) are represented as (4.34) with $r_{j,i}:=1~(i=j);~-1~(i=j+p);~0$ (otherwise) for $i\in[2p],~j\in[p]$. The formula (3.5) then implies that $\mu_{j}(\bm{\theta}^{(t-1)})=m_{j}\theta_{j}^{(t-1)}-\theta_{j}^{(t-1)}(1-\theta_{j}^{(t-1)})\frac{\partial}{\partial\theta_{j}}f(\bm{\theta}^{(t-1)}),j\in[p].$ To see (4.35), observe that for the $i$-th constraint $g_{i}(\bm{\theta})=-\theta_{i},~r_{i,i}=1,~i\in[p]$, we have
which converges to $\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{*})\geq 0$ as $\theta^{(t)}_{i}\to 0$ because it would contradict to the monotonicity of $\{f(\bm{\theta}^{(t)})\}$ otherwise. The fourth equality follows from the update formula (4.19), which implies that $m_{i}\big(\theta^{(t)}_{i}-\theta^{(t-1)}_{i}\big)=\theta^{(t)}_{i}(1-\theta^{(t-1)}_{i})\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{(t-1)})$ . Similarly, for the $(p+i)$-th constraint $g_{i+p}(\bm{\theta})=\theta_{i}-1,~r_{i+p,i+p}=-1,~i\in[p]$, we have $\nu^{(t)}_{i}=-\theta^{(t)}_{i}\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{(t)}),$ which converges to $\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{*})\geq 0$ as $\theta^{(t)}_{i}\to 1$ because it would contradict to the monotonicity of $\{f(\bm{\theta}^{(t)})\}$ otherwise. Consequently, in either constraint, we can see that $\lim\limits_{t\to\infty}\nu_{i}^{(t)}\geq 0$.
Case (ii): Algorithm 4
The simplex constraint of (4.28) corresponds to $g_{j}(\bm{\theta})=-\theta_{j},~j\in[p-1]$, and $g_{p}(\bm{\theta})=\bm{1}^{\top}\bm{\theta}-1$. With $r_{j,i}:=1~(i=j);~-1~(i=p);~0$ (otherwise) for $i\in[p],~j\in[p-1]$, the parameters $\bm{\eta}(\bm{\theta})=\big(\ln\frac{\theta_{1}}{1-\bm{1}^{\top}\bm{\theta}},\ldots,\ln\frac{\theta_{p-1}}{1-\bm{1}^{\top}\bm{\theta}}\big)^{\top}$ for the multinomial distribution are represented in the form of (4.34). The formula (3.5) then implies that
$\mu_{j}(\bm{\theta}^{(t-1)})=m\theta_{j}^{(t-1)}-\big(\theta_{j}^{(t-1)}\frac{\partial}{\partial\theta_{j}}f(\bm{\theta}^{(t-1)})-\theta_{j}^{(t-1)}(\bm{\theta}^{(t-1)})^{\top}\nabla f(\bm{\theta}^{(t-1)})\big),~j\in[p-1].$
To see (4.35), observe that for the $i$-th constraint $g_{i}(\bm{\theta})=-\theta_{i},~r_{i,i}=1,~i\in[p-1]$, we have
$\nu^{(t)}_{i}=\frac{\mu_{i}(\bm{\theta}^{(t)})-\mu_{i}(\bm{\theta}^{(t-1)})}{-\theta_{i}^{(t)}}=(1-\theta_{i}^{(t)})\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{(t)})-\sum_{j\neq i}\theta_{j}^{(t)}\frac{\partial}{\partial\theta_{j}}f(\bm{\theta}^{(t)})$ (A.10)
The second equality follows from the update formula (4.30), which implies that
$m\big(\theta^{(t)}_{i}-\theta^{(t-1)}_{i}\big)=-\theta_{i}^{(t-1)}\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{(t-1)})+\theta_{i}^{(t-1)}(\bm{\theta}^{(t-1)})^{\top}\nabla f(\bm{\theta}^{(t-1)}).$
We see from (A.10) that $\nu^{(t)}_{i}$ converges to a finite value $\nu^{*}_{i}$. To see the sign of $\nu^{*}$, observe that if $\lim\limits_{t\to\infty}\theta^{(t)}_{i}=0$, the first term of (A.10) is nonnegative because $\frac{\partial}{\partial\theta_{i}}f(\bm{\theta})$ is continuous and the optimality condition implies $\lim\limits_{t\to\infty}\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{(t)})\geq 0$. We can also show $\lim\limits_{t\to\infty}\theta^{(t)}_{i}=1$ similarly. As for the case where $\lim\limits_{t\to\infty}\theta^{(t)}_{i}\in(0,1)$, it converges to zero since $\lim\limits_{t\to\infty}\frac{\partial}{\partial\theta_{i}}f(\bm{\theta}^{(t)})=0$ must hold then. As for the second term of (A.10), we see that $\lim\limits_{t\to\infty}-\theta_{j}^{(t)}\frac{\partial}{\partial\theta_{j}}f(\bm{\theta}^{(t)})\geq 0$ for every $j\neq i$, likewise. Similarly, for the $p$-th constraint $g_{p}(\bm{\theta})=\bm{1}^{\top}\bm{\theta}-1$, we have
A.13. Proof of Proposition 5.2
@@
A.14. Proof of Proposition 5.3
@@
Let $\phi(t)=t\ln t-t+1$ and $H(\bm{x},\bm{y}):=\sum_{j\in[p]}y_{j}\phi(\frac{x_{j}}{y_{j}})=x_{j}\ln\frac{x_{j}}{y_{j}}+y_{j}-x_{j}$, and $L(\bm{\theta}):=(l_{1}(\bm{\theta}),...,l_{p}(\bm{\theta}))$ where $l_{j}(\bm{\theta}):=c_{j}-\bm{a}_{j}^{\top}\bm{\theta},j\in[p]$. Here, $H(\bm{x},\bm{y})$ denotes the KL divergence for $\bm{x}\in\mathbb{R}^{p}_{+}$ relative to $\bm{y}\in\mathbb{R}^{p}_{++}$. Observe that Algorithm 7 is a proximal-like algorithm which iteratively solves
@@
By substituting $\bm{\theta}=\bm{\theta}^{*}$, the right-hand side goes to 0 as $t\to\infty$. The second statement can be shown in the same manner as Theorem 4.3(iii) of [35] by extending $H(L(\bm{\theta}),L(\bm{\theta}^{(t)}))$ to the composite proximal term, $H(L(\bm{\theta}),L(\bm{\theta}^{(t)}))+\frac{1}{2}\|\bm{\theta}-\bm{\theta}^{(t)}\|_{\bm{\Sigma}^{-1}-\bm{Q}}^{2}$.